Write an equation of the line containing the given point and perpendicular to the given point. (0,7) 8x+3y=7
Answer by guitarheroguy(26)

Okay, first we need to find the equation of the line of 8x+3y=7, which is in standard form.
First, get the x variable on the right and change the sign
3y=-8x+7
Then divide all terms by 3
y=-8/3x+7/3
Now, perpendicular lines have the reciprical slope and different sign, plus the same y intercept
The point slope formula is y-y1=m(x-x1)
basically, all this is is y (the variable)+ the y axis coordinate (in this case 7). Equalls the slope times x minus the x axis coordinate (in this case 0)
First get the perpendicular equation of the line:
y=3/8x+7/3
Now plug in all numbers:
y-7=3/8(x-0)
Solve.
y-7=3/8x-0
Answer:
y=3/8x+7